# Dynamic Open-Vocabulary 3D Scene Graphs for Long-term Language-Guided Mobile Manipulation

### v3

@@ I Introduction
Mobile robot manipulation in real-world environments is increasingly expected to perform complex, long-term tasks that require adaptability and resilience to changes. These dynamic environments are characterized by frequent alterations caused by human activities, robot interactions, or the inherent variability of the surroundings. Traditional robotic systems often fall short in such settings because they typically rely on the assumption of a static or minimally changing environment [1, 2, 3, 4]. This limitation restricts their applicability in real-world scenarios where adaptability is crucial.
@@
Our task planning module uses the advanced large language models to decompose tasks into manageable subtasks. Then, the navigation and manipulation module are activated to execute the planned actions. In experiments, we deployed a mobile robot platform in multiple real-world indoor environments, conducting long-term task trials under varying degrees of manual environmental modifications. We demonstrated that DovSG can accurately update the 3D scene graph and performs excellent in long-term tasks and subtasks such as pick-up, place, and navigation. Our contributions are as follows:
@@ II Related Works
3D scene representation in robotics often follows two main approaches: using foundation models to create 3D structures [9, 10, 11], or combining 2D image and vision-language models [6, 8, 7, 5] to connect with the 3D world, showing impressive results in open-vocabulary tasks [12, 13]. enabling language-guided object grounding [11, 3, 12] and 3D reasoning [14]. However, dense semantic features for each point are redundant, memory-intensive, and not easily decomposable, limiting their use in dynamic robotics applications.
@@ II-3 Large Language Models for Planning
Large language models (LLMs) [17, 18] and VLMs demonstrate significant potential for zero-shot planning in robotics. These models have been used to generate trajectories and plan manipulations, improving robot adaptability. Integrating LLMs with 3D scene graphs offers further opportunities for task planning. In DovSG, we leverage GPT-4 to decompose tasks into subtasks that can be executed through graph memory, enabling robots to handle complex tasks flexibly and adaptively.
@@ III-G Manipulation
Once the robot arrives at the target location, we employ the method described in Ok-Robot [3] to perform the pick-and-place manipulation. At first, the robot’s camera is oriented toward the 3D location of the object, as provided by the semantic memory, to capture RGB-D images.
@@ IV-A Robot Setups
IV-B Task and Environment Setups
To verify whether our method enables a robot to complete long-term tasks in dynamic environments by continuously updating the scene graph, we designed an innovative experiment based on two real-world rooms. The experiment involved human interactions that modified object positions, added new objects, or revealed hidden ones, simulating a dynamic environment. Each experiment begins with a 3D scene graph as the foundation and includes two consecutive tasks. Before the robot executes the first task, we manually adjusted the scene by changing object positions, adding objects, or revealing hidden ones related to the second task, ensuring these changes could be detected during the execution of the first task.
After completing a task, instead of being repositioned to an initial start point, the robot continues to the next task from its current state, simulating real-world scenarios with continuous environmental changes caused by human interference. This approach ensures the robot not only completes individual tasks but also updates the scene graph in response to changes, maintaining an accurate understanding of the environment and efficiently completing subsequent tasks.
To evaluate the robot’s ability to detect and adapt to environmental changes, we categorized the modifications to the objects in the second task into three levels, ranging from subtle to significant changes(as shown in Fig. 4): (1) Minor Adjustment: A slight modification of the object’s position that remains detectable from the object’s original location. For instance, if the robot searches for an object in its original position, it may still be visible within the robot’s field of view, even if it’s been slightly moved. (2) Appearance: An object that was previously hidden or did not exist in the scene becomes visible due to human interaction. This could involve opening a closed drawer to reveal items inside, or a human placing a key on the table that was not there before. This change introduces new nodes to the 3D scene graph. (3) Positional Shift: A modification where the object’s location changes significantly enough that it can no longer be detected from its original position. This could involve relocating the object from one surface to another, altering its relationship with surrounding objects in the scene graph. Each task is evaluated across 40 trials, with 20 trials per room, where objects or their positions are randomized in every trial.
@@ IV-B Task and Environment Setups
SCDA 40.6% 95.3% 63.9% 92.0% 65.9% 94.3%
SGA 54.6% 89.5% 51.1% 85.2% 46.5% 82.9%
IV-C Baselines
For long-term task execution, we implemented Ok-Robot [3] on our mobile robot as a comparison method. In addition, to demonstrate our approach’s adaptability to changing environments and the effectiveness of scene graph prediction, we also evaluated scene changes and scene graph predictions. Using GPT-4o as the base model, enhanced with chain-of-thought (CoT) reasoning [28], similar to the method proposed by Jiang et al. [13], the model takes RGB observations from the robot, matches them to the most similar image in memory, detects differences, and predicts the scene graph based on the new RGB observation.
@@ IV-C Baselines
Pick up Ok-Robot 39 / 51 35 / 52 36 / 49 72.3%
Ours 53 / 68 54 / 68 53 / 66 79.2%
Place Ok-Robot 24 / 30 22 / 28 26 / 32 80.0%
Ours 33 / 41 34 / 42 33 / 39 81.9%
Navigation Ok-Robot 89 / 103 82 / 100 81 / 101 82.8%
Ours 121 / 122 118 / 124 111 / 122 95.1%
Long-term Ok-Robot 6 / 40 0 / 40 0 / 40 5.0%
Ours 16 / 40 14 / 40 10 / 40 33.3%
IV-D Evaluation
@@
To thoroughly evaluate DovSG’s adaptability to dynamic environments and its effectiveness in long-term tasks, we involved human evaluators to modify the environment, construct ground truth scene graphs, and determine task execution success. Our evaluation focuses on three key aspects: Dynamic Scene Adaptation and Scene Graph Generation: We recorded RGB observation data during the robot’s task execution and compared DovSG’s performance with baseline methods in detecting scene changes and generating scene graphs. Long-Term Tasks Evaluation: In addition to assessing the overall success rate of long-term tasks, we evaluated the success rates of individual subtasks such as picking, placing, and navigation. This detailed evaluation provides a comprehensive analysis of the effectiveness of our method across different components of task execution. The three main metrics used for evaluation are as follows: (1) Scene Change Detection Accuracy (SCDA): This metric measures the accuracy of detecting changes in the observed scene and matching them to the corresponding memory-stored RGB information. (2) Scene Graph Accuracy (SGA): This metric evaluates whether the generated scene graph matches the ground truth (GT) scene graph. If the final scene graph perfectly matches the GT version, the experiment is considered successful and assigned a score of 1, otherwise 0. (3) Task Success Rate (TSR): This metric represents the overall task completion success rate. For long-term tasks, success is only counted if all subtasks are completed successfully.
IV-E Results
@@
As shown in Tab. I, in the SCDA evaluation, GPT-4o often struggles with inconsistencies between the scope of historical observations and new observations, which introduces a significant amount of noise in the responses, resulting in subpar performance. This inconsistency also makes accurate detection under Minor Adjustment scenarios very challenging. In contrast, DovSG, supported by precise re-localization, can accurately identify the voxel index where changes have occurred in the scene, significantly outperforming the baseline. In the SGA evaluation, both approaches generate scene graphs based on new observations and identify only three types of relationships: “on”, “belong”, and “inside”. As a result, the performance in scenarios involving these three types of modifications is quite similar. The main issue with our method arises when the same object is broken into multiple nodes, which affects the accuracy of the scene graph. On the other hand, GPT-4o struggles with recognizing relationships accurately or redundantly defining objects, leading to erroneous results.
@@
In Tab. II, we have the fellow observations: (1) In the Minor Adjustment environment, despite the object being slightly moved, it remains within the robot’s field of view, which allows a high likelihood of navigating near the target. Therefore, the success rate is significantly higher compared to Positional Shift and Appearance. In the case of Positional Shift, the residual effect of CLIP features can sometimes cause the robot to mistakenly perceive the object’s location, leading it to navigate to the historical position and ultimately resulting in navigation failure. On the other hand, for Appearance, where new objects appear, the robot does not face the challenge of misjudging the original object’s location, which generally leads to a higher success rate compared to Positional Shift. (2) Compared to Ok-Robot, DovSG significantly improved the “Pick up” success rate by employing two proposed grasp strategies, focusing specifically on the target object. Additionally, we optimized the selection of grasp candidates to reduce environmental interference, ensuring that the robot chooses the optimal grasp. As a result, our method notably outperforms Ok-Robot in terms of pick-up success. (3) For long-term tasks, integrating the dynamic scene graph enables our method to significantly outperform the static-scene assumption used by Ok-Robot. Although Ok-Robot might occasionally succeed in locating the correct object in environments with minor changes, such as Minor Adjustment, it struggles with more substantial changes, such as Positional Shift or Appearance, due to its inability to perform partial updates to the scene, rendering success nearly impossible in these scenarios.
Tab. III further demonstrates the memory consumption and time required for both recording a scene and performing an update at a resolution of 1 cm for DovSG and Ok-Robot. Notably, because Ok-Robot cannot perform partial updates, the table only presents the time needed for a full scene update. The results show that Ok-Robot’s memory and time requirements are at least 20 times greater than those of DovSG.
@@
Method Resolution (cm) Memory (GB) Time (min)
Ok-Robot 1 2 20
Ours 1 0.15 1
@@ V Conclusion
In this paper, we introduce DovSG, an innovative framework designed to enable mobile robots to perform long-term tasks in dynamic environments by continuously updating 3D scene graphs. Our approach enables the robot to adapt to changes caused by human interaction or by the robot’s own task execution. Through extensive experiments in indoor environments with varying degrees of modifications, we demonstrate that DovSG can effectively detect changes in the visible scene and accurately update the local 3D scene graph. This ensures the precise execution of long-term tasks without being affected by cumulative temporal errors. By evaluating long-term tasks and subtasks such as pickup, place, and navigation, our results highlight the robustness and effectiveness of DovSG in handling the complexity of dynamic real-world scenes.

### v4

@@ I Introduction
Mobile robot manipulation in real-world environments is increasingly expected to perform complex, long-term tasks that require adaptability and resilience to changes. These dynamic environments are characterized by frequent alterations caused by human activities, robot interactions, or the inherent variability of the surroundings. Traditional robotic systems often fall short in such settings because they typically rely on the assumption of a static or minimally changing environment [1, 2, 3, 4, 5]. This limitation restricts their applicability in real-world scenarios where adaptability is crucial.
@@
Our task planning module uses the advanced large language models to decompose tasks into manageable subtasks. Then, the navigation and manipulation module are activated to execute the planned actions. In experiments, we deployed a mobile robot platform in multiple real-world indoor environments, conducting long-term task trials under varying degrees of manual environmental modifications. We demonstrated that DovSG can accurately update the 3D scene graph and perform excellently in long-term tasks and subtasks such as pick-up, place, and navigation. Our contributions are as follows:
@@ II Related Works
3D scene representation in robotics often follows two main approaches: using foundation models to create 3D structures [10, 11, 12], or combining 2D image and vision-language models [7, 9, 8, 6] to connect with the 3D world, showing impressive results in open-vocabulary tasks [13, 14], enabling language-guided object grounding [12, 3, 13] and 3D reasoning [15, 16, 17]. However, dense semantic features for each point are redundant, memory-intensive, and not easily decomposable, limiting their use in dynamic robotics applications.
@@ II-3 Large Language Models for Planning
Large language models (LLMs) [20, 21] and VLMs demonstrate significant potential for zero-shot planning in robotics [22, 23]. These models have been used to generate trajectories and plan manipulations, improving robot adaptability. Integrating LLMs with 3D scene graphs offers further opportunities for task planning. In DovSG, we leverage GPT-4 to decompose tasks into subtasks that can be executed through graph memory, enabling robots to handle complex tasks flexibly and adaptively.
@@ III-G Manipulation
Once the robot reaches the target location, we employ a pick-and-place strategy similar to Ok-Robot [3]. First, the robot retrieves the object’s 3D coordinates from semantic memory and points its camera at the target to capture RGBD images. Then, the robot performs either a “Pick up” or “Place”.
@@ IV-1 Robot Setups
IV-2 Environment and Task Setups
To verify our method’s ability to enable robots to perform long-term tasks in dynamic environments, we designed an experiment in 4 real-world rooms. The experiment simulated dynamic environments through human interactions, which modified object positions, added new objects, or revealed hidden ones. Each experiment began with a 3D scene graph and involved two consecutive tasks. Before the first task, we manually adjusted objects related to the second task by changing their positions, adding new objects, or revealing hidden ones. These modifications were designed to be detectable during the execution of the first task. After completing the first task, instead of being repositioned to an initial start point, the robot continued to the next task from its current state. This setup simulated real-world scenarios with continuous environmental changes caused by human interference. To evaluate the robot’s ability to detect and adapt to environmental changes, we categorized the modifications to the objects in the second task into three levels (see Fig. 4):
(1) Minor Adjustment: Slight position changes that remain detectable from the original location. For instance, a slightly moved object may still be visible within the robot’s field of view. (2) Appearance: Previously hidden objects become visible, such as items revealed by opening a drawer or placing a key on the table. These changes introduce new nodes into the scene graph. (3) Positional Shift: Significant relocations of objects render them undetectable from their original positions, altering their spatial relationships with other objects in the scene graph.
Overall, for each method, each level of modification was tested through 20 long-term tasks per room across 4 rooms, resulting in a total of 80 trials per modification type. With three modification levels, each method was subjected to 240 long-term task experiments, in which objects or their positions were randomized in every trial.
@@ IV-2 Environment and Task Setups
SCDA 41.44% 95.37% 64.25% 93.22% 66.35% 94.23%
SGA 54.60% 88.75% 52.25% 84.86% 46.18% 83.72%
IV-3 Baselines
To demonstrate our approach’s adaptability to changing environments and the effectiveness of scene graph prediction, we evaluated scene changes and scene graph predictions. Using GPT-4o as the base model, enhanced with chain-of-thought (CoT) reasoning [33], similar to the method proposed by Jiang et al. [14], the model takes RGB observations from the robot, matches them to the most similar image in memory, detects differences, and predicts the scene graph based on the new RGB observation. For long-term task execution, we implemented Ok-Robot [3] on our mobile robot as a comparison method. Additionally, we compared our approach with Ok-Robot and ConceptGraphs [13] in terms of memory usage and scene update time.
@@ IV-3 Baselines
Pick up Ok-Robot 84 / 110 61 / 92 59 / 87 70.58%
Ours 111 / 137 111 / 136 108 / 133 81.28%
Place Ok-Robot 53 / 64 40 / 51 42 / 51 81.32%
Ours 80 / 93 83 / 95 74 / 85 86.81%
Navigation Ok-Robot 179 / 210 146 / 184 137 / 180 80.48%
Ours 228 / 236 239 / 254 224 / 245 94.01%
Long-term Ok-Robot 12 / 80 0 / 80 0 / 80 5.00%
Ours 33 / 80 28 / 80 23 / 80 35.00%
IV-4 Evaluation
@@
To thoroughly evaluate DovSG’s adaptability to dynamic environments and its effectiveness in long-term tasks, we involved human evaluators in modifying the environment, constructing ground truth (GT) scene graphs, and determining task execution success. Our evaluation focuses on three key aspects: Dynamic Scene Adaptation and Scene Graph Generation: We recorded RGB observations during the robot’s task execution and compared DovSG’s performance with baseline methods in detecting scene changes and generating scene graphs. Long-Term Tasks Evaluation: In addition to assessing the overall success rate of long-term tasks, we evaluated the success rates of individual subtasks such as picking, placing, and navigation. This detailed evaluation provides a comprehensive analysis of the effectiveness of our method across different components of task execution. The three main metrics used for evaluation are as follows: (1) Scene Change Detection Accuracy (SCDA): This metric measures the accuracy of detecting changes in the observed scene and matching them to the corresponding memory-stored RGB information. (2) Scene Graph Accuracy (SGA): This metric evaluates whether the generated scene graph matches the GT scene graph and only recognizes three types of relations: “on”, “belong”, and “inside”. If the final scene graph perfectly matches the GT version, the experiment is considered successful and assigned a score of 1, otherwise 0. (3) Task Success Rate: This metric represents the overall task completion success rate. For long-term tasks, success is only counted if all subtasks are completed successfully.
IV-5 Results
@@
As shown in Tab. I, in the SCDA evaluation, GPT-4o often struggles with inconsistencies between the scope of historical observations and new observations, which introduces a significant amount of noise in the responses, leading to poor performance. This inconsistency also makes accurate detection under “Minor Adjustment” scenarios very challenging. In contrast, DovSG, supported by precise re-localization, can accurately identify the voxel index where changes have occurred in the scene, significantly outperforming the baseline. Specifically, in “Minor Adjustment” scenarios, DovSG exceeds the baseline by nearly 54%, primarily because GPT-4o struggles to recognize smaller positional changes. Additionally, in “Appearance” and “Positional Shift” scenarios, DovSG achieves a scene change recognition success rate approximately 28% higher than the GPT-4o. In the SGA evaluation, both approaches generate scene graphs based on new observations, with SGA accuracy in “Minor Adjustment” scenarios being higher than in “Appearance” and “Position Shift” scenarios. The main issue with our method arises when the same object is decomposed into multiple nodes, affecting the scene graph’s accuracy. On the other hand, GPT-4o struggles to accurately recognize relationships or redundantly defines objects, leading to erroneous results.
@@
In Tab. II, we have the fellow observations: (1) In the “Minor Adjustment” environment, although the object is slightly moved, it remains within the robot’s field of view. This makes it highly likely for the robot to navigate near the target, resulting in a significantly higher success rate compared to “Appearance” and “Positional Shift” (in 80 trials, it achieved 5 and 10 more successes, respectively). In the “Positional Shift” scenario, the residual effect of CLIP features can occasionally mislead the robot into navigating toward the object’s historical location, ultimately causing navigation failure. In contrast, for “Appearance”, where a new object emerges, the robot does not face the challenge of misjudging the original object’s position, generally leading to a higher success rate than “Positional Shift” (with 5 more successes out of 80 trials). (2) DovSG demonstrates significantly enhanced manipulation capabilities over Ok-Robot. For the "Pick up" task, compared to Ok-Robot, DovSG employs two proposed grasp strategies that focus specifically on the target object. By optimizing the selection of grasp candidates and integrating a heuristic grasping method, DovSG reduces environmental interference and ensures the robot selects the optimal grasp. This results in a 10.7% higher pick-up success rate than Ok-Robot, which relies solely on AnyGrasp. Regarding “Place” task, DovSG utilizes a lower placement position and an inclined placement method (see our project page for details), achieving an overall success rate that is 5.49% higher than Ok-Robot. (3) In dynamic environments, DovSG significantly outperforms Ok-Robot (which assumes a static scene) in long-term tasks, thanks to its ability to adapt to scene changes. Although Ok-Robot can occasionally succeed in locating the correct object under minor changes (e.g., “Minor Adjustment”), it struggles with larger modifications such as “Appearance” or “Positional Shift” because it cannot partially update its scene representation—making success in these scenarios nearly impossible. As a result, Ok-Robot’s success rate for long-term tasks in dynamic environments is approximately 30% lower than DovSG.
Tab. III further demonstrates the memory consumption required to store memory at a 1 cm resolution and the update times for DovSG, Ok-Robot, and ConceptGraphs after processing 1200 frames of RGBD data within a $40\,\mathrm{m}^{2}$ scene. It is important to note that since Ok-Robot and ConceptGraphs cannot perform partial updates, the table only presents the time required for a complete scene update. The results show that Ok-Robot consumes approximately 13 times more memory than DovSG and has update times 20 times longer. While ConceptGraphs and DovSG have similar memory usage due to both using 3D scene graphs, DovSG’s ability to perform local updates allows it to achieve update times 27 times faster than ConceptGraphs.
@@
Method ($\text{m}^{2}$) Memory (GB)$\downarrow$ Time (min)$\downarrow$
Ok-Robot 2 20
ConceptGraphs 0.15 27
Ours 0.15 1
@@ V Conclusion
In this paper, we introduce DovSG, an innovative framework designed to enable mobile robots to perform long-term tasks in dynamic environments by continuously local updates of 3D scene graphs. DovSG enables the robot to adapt to changes caused by human interaction or by the robot’s own task execution. These capabilities ensure the precise execution of long-term tasks without being affected by cumulative temporal errors. By evaluating long-term tasks and subtasks such as pickup, place, and navigation, our results highlight the robustness and effectiveness of DovSG in handling the complexity of dynamic real-world scenes. Compared to methods that assume static environments and cannot adjust to scene changes, DovSG achieves a 30% higher success rate in long-term tasks, accelerates updates by up to 27 times in typical room scenarios, and reduces memory consumption by a factor of 13.
Acknowledgement
This project funded by the grant from the National Natural Science Foundation of China (No. 52205244) and supported by the Ministry of Industry and Information Technology (MIIT) Key Laboratory of Intelligent Manufacturing for High-end Aerospace Products, and the Beijing Key Laboratory of Digital Design and Manufacturing.